mara is eight years older than clara. three years age, mara's age is thrice as clara. what is the present age of mara?

Answer by Theo(13342)

mara is eight years older than clara. three years age, mara's age is thrice (three times) clara's age.

what is the present age of mara?

let m = mara's age today.
let c = clara's age today.

your equations are:

m = c+8
m-3 = 3*(c-3)

in the second equation, replace m with c+8 to get c+8-3 = 3*(c-3)
simplify to get c+5 = 3c-9
subtract c from both sides of this equation and add 9 to both sides of this equation to get 14 = 2c
solve for c to get c =7

when c = 7, m = c+8 = 15
3 years ago m-3 = 12 and c-3 = 4
12 = 3*4

solution is confirmed to be good because m = c+8 and m-3 = 3*(c-3)